OpenStudy (anonymous):

On Saturday, Kelly borrowed $15 from her sister. The next day, she paid back $4. On Tuesday, she borrowed $9 more. How much money does Kelly now owe her sister?

OpenStudy (anonymous):

$20

OpenStudy (anonymous):

15 -4 + 9 = 20

OpenStudy (anonymous):

The highest and lowest temperatures ever recorded in the United States are 134 degrees at Death Valley, California, and -80 degrees at Prospect Creek, Alaska. What is the difference in these temperature extremes? @Warriorslayer

OpenStudy (anonymous):

214 degrees

OpenStudy (anonymous):

Find the difference. 8 - (-15) = @Warriorslayer

OpenStudy (anonymous):

15 + 8 = 23

OpenStudy (anonymous):

23

OpenStudy (anonymous):

A diver begins rising toward the surface from 180 feet below sea level. A few minutes later, the diver is 40 feet below sea level. How many feet did the diver rise in those few minutes?

OpenStudy (anonymous):

140

OpenStudy (anonymous):

The greenhouse effect occurs on many planets. If the actual temperature on Mars is -68 degrees, the temperature with the greenhouse effect is -63 degrees. What is the difference between the actual temperature and the temperature with the greenhouse effect?

OpenStudy (anonymous):

can anyone help me with this one?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

During the day, the temperature in Anchorage, Alaska, rose 18 degrees. The low temperature for the day was -44 degrees. What was the high temperature for the day?